10 of 2 Review Constants Periodic Part D 7.35 g of propane, C3Hg ΑΣΦ ? mol n = Submit Request Answer
Part E 260 g of Fe2O3 ? IVO AO mol n = Submit Request Answer
Answer #1

The number of moles(n) of a substance present in given mass of a sample is given by the formula:

n=\frac{mass}{molar\: \: mass}

Part D:

The molar mass of propane(C3H8) is 44.1 g/mol

The mass of propane is given to be 7.35 g

Therefore the number of moles(n) of propane present is:

n=\frac{mass}{molar\: \: mass}

n=\frac{7.35g}{44.1g/mol}=0.167\: moles

Therefore the answer is 0.167 mol

Part E:

The molar mass of Fe2O3 is 159.7 g/mol

The mass of Fe2O3 is given to be 260 g

Therefore the number of moles(n) of Fe2O3 present is:

n=\frac{mass}{molar\: \: mass}

n=\frac{260g}{159.7g/mol}=1.63\: moles

Considering the correct number of significant figures, the answer becomes 1.6 moles

Therefore the answer is 1.6 mol